# On the Completeness of Invariant Geometric Deep Learning Models

### v2

@@ E.1 Preparation
In this case, we are not given any specific mass function $m$. It seems like we need to iterate all possible mass functions to determine the number of centers, which is infeasible due to the infinite number of mass functions. However, thanks to Proposition B.5, we can determine whether $\mathcal{A}^{\text{set}}(P)$’s cardinality is 1 (i.e., whether $P$ is $\mathcal{A}$-symmetric) by checking whether the geometric centers of all corresponding sub point clouds coincide with each other.
Now, assume that there are $K$ types of node features in $X^{\mathcal{A}}$, and we use $m_{k}$ to represent the mass function that maps all $x^{\mathcal{A}}_{i}$ to 1 if $x^{\mathcal{A}}_{i}=X^{*}_{k}$ and 0 otherwise, and use $m_{c}$ to represent the mass function that maps all $x^{\mathcal{A}}_{i}$ to 1.
Based on the conclusion of center-center distance part, we know that $\{\!\!\{h_{i}^{(2)}\mid i\in[n]\}\!\!\}$ can derive the distance between $c^{m_{c}}$ to all $c^{m_{k}}$, $k\in[K]$, i.e., $\{\!\!\{h_{i}^{(2)}\mid i\in[n]\}\!\!\}\to\{d_{c^{m_{c}},c^{m_{k}}}\mid k\in[K]\}$. We can easily check the cardinality of the set $\{d_{c^{m_{c}},c^{m_{k}}}\mid k\in[K]\}$ to determine whether the geometric centers of all these $K$ sub point clouds coincide, and according to Proposition B.5, this is equivalent to determining whether $|\mathcal{A}^{\text{set}}(P)|=1$.
@@
Corollary E.5.
Given an unlabelled point cloud $P\in\mathbb{R}^{n\times 3}$. DisGNN can derive the "node-center", "center-center" and "center count" information defined in Lemma E.4 w.r.t. center sets $\mathcal{A}^{\text{set}}(P)$ after $k$ rounds, where $\mathcal{A}$ and $k$ can be:
  • $\mathcal{A}=\text{NULL}$, $k=2$. In such case, $\mathcal{A}^{\text{set}}(P)$ contains only the geometric center of $P$.
@@
In the case of where $\mathcal{A}$ is NULL, the conclusion degenerates to the Barycenter Lemma in Delle Rose et al. [10], Li et al. [37].
We now give an essential lemma. In previous work [46, 37], it has been shown that DisGNN is E(3)-incomplete, i.e., there exist pairs of non-isomorphic point clouds that DisGNN cannot distinguish. This essentially means that we cannot prove that DisGNN’s output $s$ can derive the point cloud $P$ up to permutation and Euclidean isometry. However, in the following lemma and corollary, we will show that, if we restrict the underlying space $\mathbb{P}$ to be all point clouds that are not $\mathcal{A}$-symmetric for arbitrary $\mathcal{A}$ algorithm, we can prove that DisGNN’s output can derive the input point cloud up to permutation and Euclidean isometry. This implies that DisGNN is complete on such subspaces.
Lemma E.6.
Given a point cloud $P\in\mathbb{R}^{n\times 3}_{\text{not }\mathcal{A}}$ with node features calculated by algorithm $\mathcal{A}$, $X^{\mathcal{A}}$, where $\mathbb{R}^{n\times 3}_{\text{not}\mathcal{A}}$ contains all $\mathcal{A}$-unsymmetric point clouds in $\mathbb{R}^{n\times 3}$. Then, $4$-round DisGNN’s output $s$ can derive $P$ up to permutation and Euclidean isometry.
@@
By definition of "derive", we describe the reconstruction function that takes DisGNNs’ output $s$ as input and produces the point cloud $P$ up to permutation and Euclidean isometry. We initially assume that the point cloud does not degenerate into 2D, which is a trivial case and will be discussed at the end of the proof.
Since $P$ is not $\mathcal{A}$-symmetric, there are at least two centers in $\mathcal{A}^{\text{set}}(P)$. We now take two arbitrary such centers and denote them as $c_{1}$ and $c_{2}$. According to Corollary E.5, $2$-round DisGNN’s node-level features can derive the distance information related to these two centers, namely, $h^{(2)}_{i}\to(d_{i,c_{1}},d_{i,c_{2}},d_{c_{1},c_{2}})$. In the following, we show how DisGNN can reconstruct the whole geometry based on triangular distance encoding.
@@
The coordinates of these nodes are added to the “known" set of nodes $K$, while the remaining nodes are labeled as “unknown."
It is established that there are no “unknown" nodes in the interior (as $\theta_{xc_{1}c_{2},yc_{1}c_{2}}$ is minimal) and border (since we have identified these points) of the $x$-plane and $y$-plane. We denote the interior and border of the two planes as $P_{0}$.
Now, we reflect the $y$-plane mirrorly w.r.t. the $x$-plane which yields the $p_{-1}$-plane. The interior and border formed by $p_{-1}$-plane and $x$-plane are denoted as $P_{-1}$. Since there are no “unknown" nodes in $P_{0}$, the remaining undetermined nodes in $P_{-1}$ can be determined from $h_{x}^{(4)}$, as their another possible position calculated by $h_{x}^{(4)}$ lie in the “dead area" $P_{0}$, where it is ensured to be empty (no undetermined nodes left). Similarly, reflecting the $x$-plane w.r.t. the $y$-plane produces the $p_{1}$-plane, and the corresponding area by $p_{1}$-plane and $y$-plane is denoted as $P_{1}$. Nodes in $P_{1}$ are determined likewise from $h_{y}^{(3)}$. This process continues by reflecting the $p_{1}$-plane w.r.t. the $x$-plane to obtain the $p_{-2}$-plane, determining the nodes in $P_{-2}$, and so forth, until all $P$ areas collectively cover the entire 3D space. Consequently, all nodes can be determined, and the geometric structure can be reconstructed.
@@
As an insightful takeaway from Lemma E.6, we notice that the key in reconstruction is the initial triangular distance encoding that can be captured by DisGNN. The triangular distance encoding enhances each node $i$’s feature with distance information $(d_{ic_{1}},d_{ic_{2}},d_{c_{1}c_{2}})$, where $c_{1}$ and $c_{2}$ are two global anchors. In Lemma E.6, the two anchors are from $\mathcal{A}$ center sets, and DisGNN can capture the corresponding distance encoding according to Lemma E.4. In Section 4.1, we show that we can mark a node as an additional anchor, which can also be captured by DisGNN due to the distinct mark. Other designs may also be proposed by considering this insightful idea.
Similarly to Corollary E.5, DisGNN can take unlabelled point clouds as input and finally reconstruct them when the unlabelled point clouds are $\mathcal{C}$-unsymmetric or $\mathcal{D}$-unsymmetric.
Corollary E.7.
Given a point cloud $P\in\mathbb{R}^{n\times 3}$, assume that $P$ is $\mathcal{A}$-unsymmetric. Then, $k$-round DisGNN’s output $s$ can derive $P$ up to permutation and Euclidean isometry, where $\mathcal{A}$ and $k$ can be:
@@
Theorem 3.2. (Unsymmetric point clouds are identifiable) Let $\mathcal{A}$ be either center distance encoding $\mathcal{C}$ or DisGNN encoding $\mathcal{D}$. Then, given an arbitrary point cloud $P\in\mathbb{R}^{n\times 3}$, $P$ is not $\mathcal{A}$-symmetric $\Rightarrow$ $P$ can be identified by DisGNN.
@@
It suffices to show that $P$ is not $\mathcal{D}$-symmetric $\Rightarrow$ $P$ can be identified by DisGNN by Proposition E.3.
@@
  • $P^{\prime}$ is not $\mathcal{D}$-symmetric. As a direct consequence of Lemma E.6 and Remark E.2, DisGNN can distinguish $P$ and $P^{\prime}$.
  • $P^{\prime}$ is $\mathcal{D}$-symmetric. According to Lemma E.4 and Corollary E.5, DisGNN’s output for $P$ and $P^{\prime}$ can derive $\mathds{1}_{|\mathcal{D}^{\text{set}}(P)|=1}$ and $\mathds{1}_{|\mathcal{D}^{\text{set}}(P^{\prime})|=1}$ respectively. Since $P$ is not $\mathcal{D}$-symmetric while $P^{\prime}$ is, by definition, we know that $\mathds{1}_{|\mathcal{D}^{\text{set}}(P)|=1}\neq\mathds{1}_{|\mathcal{D}^{\text{set}}(P^{\prime})|=1}$, and thus DisGNN can distinguish $P$ and $P^{\prime}$.
E.3 Proof of Theorem 4.2
@@
    After 2 rounds of DisGNN, according to Lemma E.4, for all $j\in[n]$, $h_{j}^{(2)}\to d_{jc}$.
@@
At this point, for all nodes $j$, $h_{j}^{(3)}$ can derive $d_{ji}$, $d_{jc}$, and $d_{ic}$, thus completing the triangular distance encoding necessary for reconstruction in Lemma E.6. Therefore, similar to the proof of Lemma E.6, with another two more rounds of DisGNN, the output can derive the point cloud up to permutation and Euclidean isometry.
@@
There is still a potential problem with the above proof: node $i$ may coincide with the geometric center. In such case, we can not anymore obtain triangle distance information formed by each node $j$ and $i,c$, which is essential in Lemma E.6 for reconstruction. However, since $P$ has more than 2 nodes, there is at least one node that satisfies the assumption.
Notice that GeoNGNN injectively pools all the subgraph representations in the outer GNN and obtains the final output $s=f_{\text{outer}}(\{\!\!\{s_{i}\mid i\in[n]\}\!\!\})$. Each subgraph represnetation $s_{i}$ can derive the distence between node $i$ and the geometric center $c$ as following: $s_{i}\to h_{i}^{(5)}\to h_{i}^{(4)}\to...\to h_{i}^{(2)}\to d_{ic}$, according to Lemma E.4. This can be leveraged as an indicator, telling us whether node $i$ coincides with the $c$. As a consequence, we can derive the point cloud from $s$ as follows: first, search across all subgraph representations $s_{i}$ to find the one that does not coincide with the geometric center, then reconstruct the point cloud based on this subgraph representation according to Lemma E.6. This finishes the proof.
E.4 Proof of Theorem 4.4
@@
We establish by demonstrating that all general geometric subgraph GNNs $\mathsf{GeoA}(\mathcal{A},\mathsf{Pool},L,r_{\text{sub}}=+\infty,r_{\text{cutoff}}=+\infty)$, abbreviated as $\mathsf{GeoA}$ hereafter, are capable of implementing [17] GeoNGNN (with $N_{\text{inner}}$ inner layers, $r_{\text{sub}}=+\infty,r_{\text{cutoff}}=+\infty$ and without outer GNN) – denoted as complete GeoNGNN – with a fixed number of layers $L$. Since complete GeoNGNN is E(3)-complete when $N_{\text{inner}}\geq 5$, $\mathsf{GeoA}$ can also achieve E(3)-completeness when $L$ is larger than some constant.
@@
In the scenario where $\mathsf{Geoagg}^{\mathsf{L}}_{\mathsf{u}}\in\mathcal{A}$, a single aggregation layer in $\mathsf{GeoA}$ can implement a complete GeoNGNN layer by learning an aggregation function $f^{(l)}$, defined in Section D, that only maintains $\mathsf{Geoagg}^{\mathsf{P}}_{\mathsf{u,v}}$ and $\mathsf{Geoagg}^{\mathsf{L}}_{\mathsf{u}}$ while disregarding other aggregation operations.
@@
$\to$ $\{\!\!\{h^{(L_{\text{SV}})}(v,v)\mid v\in\mathcal{V}_{G}\}\!\!\}$
$\to$ $\{\!\!\{\{\!\!\{h^{(L_{\text{VS}})}(v,u)\mid u\in\mathcal{V}_{G}\}\!\!\}\mid v\in\mathcal{V}_{G}\}\!\!\}$
@@
Proposition E.8.
Let $\mathsf{GeoA}$ be any general geometric subgraph GNN defined in Section D. Denote $\mathcal{A}^{\mathsf{u}\leftrightarrow\mathsf{v}}$ as the aggregation scheme obtained from $\mathcal{A}$ by exchanging the element $\mathsf{agg}^{\mathsf{P}}_{\mathsf{uu}}$ with $\mathsf{agg}^{\mathsf{P}}_{\mathsf{vv}}$, exchanging $\mathsf{Geoagg}^{\mathsf{L}}_{\mathsf{u}}$ with $\mathsf{Geoagg}^{\mathsf{L}}_{\mathsf{v}}$, and exchanging $\mathsf{agg}^{\mathsf{G}}_{\mathsf{u}}$ with $\mathsf{agg}^{\mathsf{G}}_{\mathsf{v}}$. Then, $\mathsf{GeoA}(\mathcal{A},\mathsf{VS},L,r_{\text{sub}},r_{\text{cutoff}})$ and $\mathsf{GeoA}(\mathcal{A}^{\mathsf{u}\leftrightarrow\mathsf{v}},\mathsf{SV},L,r_{\text{sub}},r_{\text{cutoff}})$ can implement each other.
@@
Similar to the original proof of Proposition 4.5. in Zhang et al. [62], the proof is almost trivial by symmetry: All functions within $\mathsf{GeoA}(\mathcal{A},\mathsf{VS})$ can inherently learn to ensure that $h^{(l)}_{uv}$ in $\mathsf{GeoA}(\mathcal{A},\mathsf{VS})$ precisely corresponds to $h^{(l)}(v,u)$ in $\mathsf{GeoA}(\mathcal{A}^{\mathsf{u}\leftrightarrow\mathsf{v}},\mathsf{SV})$ for any arbitrary $u,v,l$, with the reverse equivalence also holding. Additionally, given the symmetry of the pooling method, consistency in the output can be guaranteed. ∎
As a direct consequence of Proposition E.8, all of the general subgraph GNN with $\mathsf{Geoagg}^{\mathsf{L}}_{\mathsf{v}}$ can implement another one with $\mathsf{Geoagg}^{\mathsf{L}}_{\mathsf{u}}$, and according to the previous proof, they are also E(3)-complete under the conditions specified in Theorem. This ends the proof.
E.5 Proof of Theorem 4.5
Theorem 4.5 (E(3)-Completeness of DimeNet, SphereNet, GemNet) When the following conditions are met, DimeNet, SphereNet [^5] Here in SphereNet we do not consider the relative azimuthal angle $\varphi$, since SphereNet with $\varphi$ is not E(3)-invariant, while E(3)-completeness is defined on E(3)-invariant models. Note that the exclusion of $\varphi$ only results in weaker expressiveness. and GemNet are E(3)-complete.
  • The aggregation layer number is larger than some constant $C$, which is irrelevant to the node number.
@@
The main idea to prove this theorem is to also show that DimeNet, SphereNet and GemNet can implement [17] GeoNGNN, which is E(3)-complete, and thus they are E(3)-complete as well. We elaborate on these models in separate subsections.
E.5.1 E(3)-completeness of DimeNet

### v3

@@ B.2 Preparation for Reconstruction Proof
In this case, we are not given any specific mass function $m$. It seems like we need to iterate all possible mass functions to determine the number of centers, which is infeasible due to the infinite number of mass functions. However, thanks to Proposition C.4, we can determine whether $\mathcal{A}^{\text{set}}(P)$’s cardinality is 1 (i.e., whether $P$ is $\mathcal{A}$-symmetric) by checking whether the geometric centers of all corresponding sub point clouds coincide with each other.
Now, assume that there are $K$ types of node features in $X^{\mathcal{A}}$, and we use $m_{k}$ to represent the mass function that maps all $x^{\mathcal{A}}_{i}$ to 1 if $x^{\mathcal{A}}_{i}$ is the $k$-th type of features and 0 otherwise, and use $c$ solely to represent the coordinate of geometric center (unweighted average coordinates of all points).
Based on the conclusion of center-center distance part, we know that $\{\!\!\{h_{i}^{(2)}\mid i\in[n]\}\!\!\}$ can derive the distance between $c$ to all $c^{m_{k}}$, $k\in[K]$, i.e., $\{\!\!\{h_{i}^{(2)}\mid i\in[n]\}\!\!\}\to\{d_{c,c^{m_{k}}}\mid k\in[K]\}$. We can easily check the cardinality of the set $\{d_{c,c^{m_{k}}}\mid k\in[K]\}$ to determine whether the geometric centers of all these $K$ sub point clouds coincide, and according to Proposition C.4, this is equivalent to determining whether $|\mathcal{A}^{\text{set}}(P)|=1$.
@@
Corollary B.5.
Given an unlabelled point cloud $P\in\mathbb{R}^{n\times 3}$. DisGNN can derive the “node-center”, “center-center” and “center count” information defined in Lemma B.4 w.r.t. center sets $\mathcal{A}^{\text{set}}(P)$ after $k$ rounds, where $\mathcal{A}$ and $k$ can be:
  • $\mathcal{A}=\text{NULL}$ (i.e., $\mathcal{A}$ assigns the same features to all nodes), $k=2$. In such case, $\mathcal{A}^{\text{set}}(P)$ contains only the geometric center of $P$.
@@
In the case of where $\mathcal{A}$ is NULL, the conclusion degenerates to the Barycenter Lemma in Delle Rose et al. (2023); Li et al. (2024).
We now give an essential lemma. In previous work (Pozdnyakov & Ceriotti, 2022; Li et al., 2024), it has been shown that DisGNN is E(3)-incomplete, i.e., there exist pairs of non-isomorphic point clouds that DisGNN cannot distinguish. This essentially means that we cannot prove that DisGNN’s output $s$ can derive the point cloud $P$ up to permutation and Euclidean isometry. However, in the following lemma and corollary, we will show that, if we restrict the underlying space $\mathbb{P}$ to be all point clouds that are not $\mathcal{A}$-symmetric for arbitrary $\mathcal{A}$ algorithm, we can prove that DisGNN’s output can derive the input point cloud up to permutation and Euclidean isometry. This implies that DisGNN is complete on such subsets.
Lemma B.6.
Given a point cloud $P\in\mathbb{R}^{n\times 3}_{\text{not }\mathcal{A}}$ with node features $X^{\mathcal{A}}$ calculated by algorithm $\mathcal{A}$, where $\mathbb{R}^{n\times 3}_{\text{not }\mathcal{A}}$ contains all $\mathcal{A}$-asymmetric point clouds in $\mathbb{R}^{n\times 3}$. Then, $4$-round DisGNN’s output $s$ can derive $P$ up to permutation and Euclidean isometry.
@@
By definition of “derive”, we describe the reconstruction function that takes DisGNNs’ output $s$ as input and produces the point cloud $P$ up to permutation and Euclidean isometry. We initially assume that the point cloud does not degenerate into 2D, which is a trivial case and will be discussed at the end of the proof.
Since $P$ is $\mathcal{A}$-asymmetric, there are at least two centers in $\mathcal{A}^{\text{set}}(P)$. We now take two arbitrary such centers and denote them as $c_{1}$ and $c_{2}$. According to Corollary B.5, $2$-round DisGNN’s node-level features can derive the distance information related to these two centers, namely, $h^{(2)}_{i}\to(d_{i,c_{1}},d_{i,c_{2}},d_{c_{1},c_{2}})$. In the following, we show how DisGNN can reconstruct the whole geometry based on triangular distance encoding.
@@
The coordinates of these nodes are added to the “known” set of nodes $K$, while the remaining nodes are labeled as “unknown.”
It is established that there are no “unknown” nodes in the interior (as $\theta_{xc_{1}c_{2},yc_{1}c_{2}}$ is minimal) and border (since we have identified these points) of the $x$-plane and $y$-plane. We denote the interior and border of the two planes as $P_{0}$.
Now, we reflect the $y$-plane mirrorly w.r.t. the $x$-plane which yields the $p_{-1}$-plane. The interior and border formed by $p_{-1}$-plane and $x$-plane are denoted as $P_{-1}$. Since there are no “unknown” nodes in $P_{0}$, the remaining undetermined nodes in $P_{-1}$ can be determined from $h_{x}^{(4)}$, as their another possible position calculated by $h_{x}^{(4)}$ lie in the “dead area” $P_{0}$, where it is ensured to be empty (no undetermined nodes left). Similarly, reflecting the $x$-plane w.r.t. the $y$-plane produces the $p_{1}$-plane, and the corresponding area by $p_{1}$-plane and $y$-plane is denoted as $P_{1}$. Nodes in $P_{1}$ are determined likewise from $h_{y}^{(3)}$. This process continues by reflecting the $p_{1}$-plane w.r.t. the $x$-plane to obtain the $p_{-2}$-plane, determining the nodes in $P_{-2}$, and so forth, until all $P$ areas collectively cover the entire 3D space. Consequently, all nodes can be determined, and the geometric structure can be reconstructed.
@@
As an insightful takeaway from Lemma B.6, we notice that the key in reconstruction is the initial triangular distance encoding that can be captured by DisGNN. The triangular distance encoding enhances each node $i$’s feature with distance information $(d_{ic_{1}},d_{ic_{2}},d_{c_{1}c_{2}})$, where $c_{1}$ and $c_{2}$ are two global anchors. In Lemma B.6, the two anchors are from $\mathcal{A}$ center sets, and DisGNN can capture the corresponding distance encoding according to Lemma B.4. In Section 5.1, we show that we can mark a node as an additional anchor, which can also be captured by DisGNN due to the distinct mark. Other designs may also be proposed by considering this insightful idea.
Similarly to Corollary B.5, DisGNN can take unlabelled point clouds as input and finally reconstruct them when the unlabelled point clouds are $\mathcal{C}$-asymmetric or $\mathcal{D}$-asymmetric.
Corollary B.7.
Given a point cloud $P\in\mathbb{R}^{n\times 3}$, assume that $P$ is $\mathcal{A}$-asymmetric. Then, $k$-round DisGNN’s output $s$ can derive $P$ up to permutation and Euclidean isometry, where $\mathcal{A}$ and $k$ can be:
@@
Theorem 4.2. (Asymmetric point clouds are identifiable) Let $\mathcal{C}$ denote the center distance encoding and $\mathcal{D}$ the DisGNN encoding. Then, given an arbitrary point cloud $P\in\mathbb{R}^{n\times 3}$, $P$ is $\mathcal{C}$-asymmetric $\Rightarrow$ $P$ is $\mathcal{D}$-asymmetric $\Rightarrow$ $P$ can be identified by DisGNN.
@@
The first $\Rightarrow$ is proved in Proposition B.3.
It suffices to show that $P$ is $\mathcal{D}$-asymmetric $\Rightarrow$ $P$ can be identified by DisGNN.
@@
  • $P^{\prime}$ is $\mathcal{D}$-asymmetric. As a direct consequence of Lemma B.6Corollary B.7 and Proposition B.2, DisGNN can distinguish $P$ and $P^{\prime}$.
  • $P^{\prime}$ is $\mathcal{D}$-symmetric. According to Lemma B.4 and Corollary B.5, DisGNN’s output for $P$ and $P^{\prime}$ can derive $\mathds{1}_{|\mathcal{D}^{\text{set}}(P)|=1}$ and $\mathds{1}_{|\mathcal{D}^{\text{set}}(P^{\prime})|=1}$ respectively. Since $P$ is $\mathcal{D}$-asymmetric while $P^{\prime}$ is, by definition, we know that $\mathds{1}_{|\mathcal{D}^{\text{set}}(P)|=1}\neq\mathds{1}_{|\mathcal{D}^{\text{set}}(P^{\prime})|=1}$, and thus DisGNN can distinguish $P$ and $P^{\prime}$.
B.4 Proof of Theorem 5.1
@@
    After 2 rounds of DisGNN, according to Lemma B.4, for all $j\in[n]$, $h_{j}^{(2)}\to d_{jc}$.
@@
At this point, for all nodes $j$, $h_{j}^{(3)}$ can derive $d_{ji}$, $d_{jc}$, and $d_{ic}$, thus completing the triangular distance encoding necessary for reconstruction in Lemma B.6. Therefore, similar to the proof of Lemma B.6, with another two more rounds of DisGNN, the output can derive the point cloud up to permutation and Euclidean isometry.
@@
There is still a potential problem with the above proof: node $i$ may coincide with the geometric center. In such case, we can not anymore obtain triangle distance information formed by each node $j$ and $i,c$, which is essential in Lemma B.6 for reconstruction. However, since $P$ has more than 2 nodes, there is at least one node that satisfies the assumption.
Notice that GeoNGNN injectively pools all the subgraph representations in the outer GNN and obtains the final output $s=f_{\text{outer}}(\{\!\!\{s_{i}\mid i\in[n]\}\!\!\})$. Each subgraph represnetation $s_{i}$ can derive the distence between node $i$ and the geometric center $c$ as following: $s_{i}\to h_{i}^{(5)}\to h_{i}^{(4)}\to...\to h_{i}^{(2)}\to d_{ic}$, according to Lemma B.4. This can be leveraged as an indicator, telling us whether node $i$ coincides with the $c$. As a consequence, we can derive the point cloud from $s$ as follows: first, search across all subgraph representations $s_{i}$ to find the one that does not coincide with the geometric center, then reconstruct the point cloud based on this subgraph representation according to Lemma B.6. This finishes the proof.
B.5 Proof of Theorem 5.3
@@
We establish by demonstrating that all general geometric subgraph GNNs $\mathsf{GeoA}({\mathcal{A}},\mathsf{Pool},L,r_{\text{sub}}=+\infty,r_{\text{cutoff}}=+\infty)$ (Please see Appendix E for formal definitions), abbreviated as $\mathsf{GeoA}$ hereafter, are capable of implementing (Frasca et al., 2022) GeoNGNN (with $N_{\text{inner}}$ inner layers, $r_{\text{sub}}=+\infty,r_{\text{cutoff}}=+\infty$ and without outer GNN) – denoted as complete GeoNGNN – with a fixed number of layers $L$. Since complete GeoNGNN is E(3)-complete when $N_{\text{inner}}\geq 5$, $\mathsf{GeoA}$ can also achieve E(3)-completeness when $L$ is larger than some constant.
@@
In the scenario where $\mathsf{Geoagg}^{\mathsf{L}}_{\mathsf{u}}\in{\mathcal{A}}$, a single aggregation layer in $\mathsf{GeoA}$ can implement a complete GeoNGNN layer by learning an aggregation function $f^{(l)}$, defined in Section E, that only maintains $\mathsf{Geoagg}^{\mathsf{P}}_{\mathsf{u,v}}$ and $\mathsf{Geoagg}^{\mathsf{L}}_{\mathsf{u}}$ while disregarding other aggregation operations.
@@
$\to$ $\{\!\!\{h^{(L_{\text{SV}})}_{vv}\mid v\in\mathcal{V}_{G}\}\!\!\}$
$\to$ $\{\!\!\{\{\!\!\{h^{(L_{\text{VS}})}_{vu}\mid u\in\mathcal{V}_{G}\}\!\!\}\mid v\in\mathcal{V}_{G}\}\!\!\}$
@@
Proposition B.8.
Let $\mathsf{GeoA}$ be any general geometric subgraph GNN defined in Section E. Denote ${\mathcal{A}}^{\mathsf{u}\leftrightarrow\mathsf{v}}$ as the aggregation scheme obtained from ${\mathcal{A}}$ by exchanging the element $\mathsf{agg}^{\mathsf{P}}_{\mathsf{uu}}$ with $\mathsf{agg}^{\mathsf{P}}_{\mathsf{vv}}$, exchanging $\mathsf{Geoagg}^{\mathsf{L}}_{\mathsf{u}}$ with $\mathsf{Geoagg}^{\mathsf{L}}_{\mathsf{v}}$, and exchanging $\mathsf{agg}^{\mathsf{G}}_{\mathsf{u}}$ with $\mathsf{agg}^{\mathsf{G}}_{\mathsf{v}}$. Then, $\mathsf{GeoA}({\mathcal{A}},\mathsf{VS},L,r_{\text{sub}},r_{\text{cutoff}})$ and $\mathsf{GeoA}({\mathcal{A}}^{\mathsf{u}\leftrightarrow\mathsf{v}},\mathsf{SV},L,r_{\text{sub}},r_{\text{cutoff}})$ can implement each other.
@@
Similar to the original proof of Proposition 4.5. in Zhang et al. (2023), the proof is almost trivial by symmetry: All functions within $\mathsf{GeoA}({\mathcal{A}},\mathsf{VS})$ can inherently learn to ensure that $h^{(l)}_{uv}$ in $\mathsf{GeoA}({\mathcal{A}},\mathsf{VS})$ precisely corresponds to $h^{(l)}_{vu}$ in $\mathsf{GeoA}({\mathcal{A}}^{\mathsf{u}\leftrightarrow\mathsf{v}},\mathsf{SV})$ for any arbitrary $u,v,l$, with the reverse equivalence also holding. Additionally, given the symmetry of the pooling method, consistency in the output can be guaranteed. ∎
As a direct consequence of Proposition B.8, all of the general subgraph GNN with $\mathsf{Geoagg}^{\mathsf{L}}_{\mathsf{v}}$ can implement another equivalent one with $\mathsf{Geoagg}^{\mathsf{L}}_{\mathsf{u}}$, and according to the previous proof, they are also E(3)-complete under the conditions specified in Theorem. This ends the proof.
B.6 Proof of Theorem 5.4
Theorem 5.4 (E(3)-Completeness of DimeNet, SphereNet, GemNet) When the following conditions are met, DimeNet, SphereNet and GemNet are E(3)-complete.
  • The aggregation layer number is larger than some constant $C$ (irrelevant to the node number).
@@
  • They interact with all neighbors, i.e., $r_{\text{int}}=+\infty$..
The main idea to prove this theorem is to also show that DimeNet, SphereNet and GemNet can implement (Frasca et al., 2022) GeoNGNN, which is E(3)-complete, and thus they are E(3)-complete as well. We elaborate on these models in separate subsections.
B.6.1 E(3)-completeness of DimeNet